An educational psychologist wishes to know the mean number of words a third grader can read per minute. She wants to ensure that the estimate has an error of at most 0.25 words per minute. a previous study found that the mean was 25.8 words per minute. assuming that the variance is 9.61. what is the minimum number of third graders that must be included in a sample to construct the 98% confidence interval
Answer #1

Answer:

Given,

Margin of error E = 0.25

Mean = 25.8

Variance = 9.61

Standard deviation = sqrt(variance)

= sqrt(9.61)

= 3.1

Here for 98% CI, z value is 2.33

Sample size n = (z*s/E)^2

substitute values

= (2.33*3.1/0.25)^2

= 834.75

n = 835